### v2

# Senior Authors

Abstract
Computer-use agents face a fundamental limitation. They rely exclusively on primitive GUI actions (click, type, scroll), creating brittle execution chains prone to cascading failures. While API-driven agents harness rich capabilities through structured interfaces and tools, computer-use agents remain constrained to low-level visual interactions. We present UltraCUA, a foundation model that transcends this limitation through hybrid action—seamlessly unifying primitive GUI operations with high-level tool execution. Our innovation rests on four critical advances. First, an automated pipeline extracts and scales tool capabilities from software documentation and code repositories. Second, a synthetic data engine produces 17,000+ verifiable tasks capturing real-world computer-use complexity. Third, comprehensive hybrid action trajectory collection incorporates both GUI primitives and strategic tool calls. Fourth, a two-stage training methodology combines supervised fine-tuning with online reinforcement learning, enabling intelligent action selection between GUI and API. Evaluation with our 7B and 32B UltraCUA models reveals transformative performance gains. On OSWorld, UltraCUA achieves 22% relative improvement while executing 11% faster than existing approaches, averagely. Cross-domain validation on WindowsAgentArena demonstrates robust generalization with 21.7% success rate, surpassing Windows-trained baselines. The hybrid action paradigm proves essential, reducing error propagation while improving execution efficiency. This work establishes a scalable paradigm bridging primitive GUI interactions and high-level tool intelligence, enabling more resilient and adaptable computer use agents for diverse environments and complex real-world tasks.
(a) OSWorld and WindowsAgentArena
(b) Existing Agents v.s. UltraCUA
Figure 1: (a) UltraCUA demonstrates superior performance on OSWorld and WindowsAgentArena benchmarks; (b) Comparison of execution paradigms between existing GUI agents and UltraCUA. Traditional agents using pure low-level actions suffer from cascading errors, while UltraCUA’s hybrid approach achieves faster and more robust execution.
1 Introduction
Computer-use automation has emerged as a transformative capability for autonomous agents. It enables interaction with the vast ecosystem of desktop and web applications that define modern digital workflows (Hong et al., 2023; Shaw et al., 2024; Zhang et al., 2023). However, current computer-use agents (CUAs) face a critical architectural limitation. They operate exclusively through primitive visual actions such as clicking, typing, and scrolling (Rawles et al., 2024; Koh et al., 2024). This constraint creates a profound performance chasm compared to API-driven agents that harness rich tool-based interfaces—MCP servers and specialized tools to accomplish complex tasks with efficiency (Qin et al., 2023b; Schick et al., 2023b).
The exclusive reliance on primitive actions introduces fundamental challenges that severely limit agent effectiveness. First, extended execution chains accumulate errors that propagate into catastrophic failures. A single misplaced click can completely derail complex workflows (Zheng et al., 2024; Yan et al., 2023). Second, operations achievable through single tool calls require dozens of fragile GUI interactions. This creates severe performance bottlenecks. For instance, extracting data from multiple spreadsheets forces traditional CUAs to navigate complex menus, manually select individual cells, copy values, switch applications, and paste content. Each step represents a critical failure point. Conversely, an agent with native spreadsheet APIs could accomplish this task reliably with minimal operations. This efficiency disparity is striking. While agents leveraging tool-based interfaces consistently exceed 80% success on challenging benchmarks like GAIA (Mialon et al., 2024; Zhang et al., 2025), GUI-constrained computer-use agents remain fundamentally bottlenecked. This compels our unified approach that marries GUI universality with tool-based precision.
In this paper, we eliminate this capability gap through hybrid action, a paradigm that recognizes the complementary strengths of both execution modalities. Our approach seamlessly integrates GUI primitives with high-level tool execution, unlocking the best of both worlds: the universal applicability of visual interactions and the efficiency of direct tool access. Rather than treating these paradigms as mutually exclusive alternatives, we enable intelligent orchestration that maximizes task success while minimizing execution overhead. Agents learn to strategically deploy tool calls when they offer clear efficiency advantages, while preserving GUI interactions for comprehensive coverage and granular control when tools are unavailable or insufficient.
To summarize, our technical contributions include:
  • An automated pipeline for synthesizing tool capabilities that transcends manual curation limitations (Qin et al., 2023a; Tang et al., 2023). Unlike existing approaches that rely on fixed tool sets, our system dynamically extracts tools from software documentation, integrates open-sourced implementations, and employs coding agents to generate tools on demand. This enables scalable tool acquisition across diverse environments, producing hundreds of tools that form a critical foundation for hybrid action execution.
  • A dual-pipeline synthetic data engine that solves the fundamental challenge of verifiable task generation for computer-use agent training. Traditional approaches fail to reliably assess task completion in dynamic environments, creating unreliable training data. We develop two innovative pipelines that together produce 17,000+ tasks with guaranteed verification. The instruction-first pipeline leverages agent exploration to generate contextually relevant tasks, with dedicated evaluators ensuring execution validity. The evaluator-first pipeline reverses this process by collecting atomic verification functions from target environments, then reprogramming or composing them for complex assessment criteria. Tasks are generated to satisfy these pre-built evaluators, ensuring reliable completion detection. This breakthrough eliminates the verification bottleneck that has limited training computer-use agents at scale.
  • A large-scale hybrid action trajectory collection that bridges the critical gap between pure GUI datasets and tool-augmented execution. Existing computer-use datasets contain only primitive GUI action sequences. They lack demonstrations of strategic tool integration, severely limiting agent learning capabilities. We address this fundamental limitation by collecting 20,000+ successful trajectories using an multi-agent framework. Our approach combines a powerful planner (e.g., OpenAI o3) with a state-of-the-art grounding model (e.g., GTA1-7B (Yang et al., 2025a)). The planner intelligently selects between tool calls and GUI actions based on efficiency considerations. The grounder ensures precise execution of chosen actions. This dataset uniquely captures the decision-making patterns required for seamless action mode transitions, enabling agents to learn when and how to leverage hybrid capabilities.
  • An agentic foundation model with hybrid action intelligence that transcends the limitations of GUI-only training. We develop models at two scales (7B and 32B) through our data and tool ecosystem. Our training employs a strategic two-stage methodology. First, supervised fine-tuning on high-quality hybrid trajectories teaches agents when to leverage tools versus GUI actions. Second, online reinforcement learning on verifiable synthetic tasks refines decision-making through on-policy feedback. This approach produces agents that strategically balance tool efficiency with GUI universality.
Experiments reveal the transformative impact of UltraCUA’s hybrid action across diverse benchmarks and platforms. On OSWorld (Xie et al., 2024), our UltraCUA models deliver a compelling 22% relative improvement over base models at 7B and 32B scales, averagely. In terms of cross-platform generalization, UltraCUA-7B achieves 21.7% success on WindowsAgentArena (Bonatti et al., 2024) without any Windows-specific training, surpassing models explicitly trained on Windows data. This demonstrates that UltraCUA’s hybrid action capabilities fundamentally enhance agent adaptability rather than merely optimizing for specific environments. The consistent gains across model scales and operating systems establish UltraCUA as a new paradigm for robust computer-use automation.
2 Methodology
Figure 2: An overview of UltraCUA’s design. UltraCUA adaptively switches between visual grounding and tool execution based on task requirements. This establishes the hybrid action mechanism that seamlessly integrates GUI primitives with high-level tool capabilities for enhanced computer-use automation across diverse environments.
Our methodology comprises three key components for developing a foundation CUA model with hybrid action. First, we build a comprehensive collection of programmatic tools through an automated extraction pipeline. Second, we design a dual-pipeline synthetic data engine that generates verifiable tasks for complex real-world computer use. Finally, we train our model via supervised fine-tuning on collected trajectories followed by online reinforcement learning on synthetic tasks.
2.1 Automated Tool Collection for Hybrid Action
The foundation of our approach is hybrid action—seamlessly integrating primitive GUI actions with high-level programmatic tools. We define a “tool” as a high-level interface encapsulating sequences of computer-use actions, typically implemented as Python functions, keyboard shortcuts, or combinations of primitive actions (e.g., type, key combinations)—but excluding actions requiring visual grounding like clicks. Each tool is exposed to the model through a Python function signature with descriptive docstrings specifying parameters and functionality.
While GUI-only agents suffer from cascading failures in lengthy action sequences, programmatic interfaces alone cannot handle all computer interactions. Our hybrid approach enables agents to leverage programmatic tools for efficiency when available, while retaining GUI actions to ensure generalization. To build a hybrid action space where tools cover diverse applications and usage scenarios, we developed an automated pipeline collecting hundreds of tools from the following three complementary sources. Tool details are also present in Appendix A.3.
Extraction from Software Documentation. Application documentation contains expert knowledge, particularly keyboard shortcuts, that bypass tedious GUI sequences. For example, changing VS Code’s color theme requires navigating File → Preferences → Color Theme with GUI actions. Our pipeline extracts the shortcut (Ctrl+K, Ctrl+T) from documentation and converts it into a programmatic tool: vscode.set_theme(). This transforms fragile multi-step sequences into single, reliable operations.
Integration of Open-Source Implementations. We incorporate existing programmatic tools from open-sourced frameworks, particularly leveraging implementations from AgentS2 (Agashe et al., 2025) and AgentStore (Jia et al., 2024). These tools transform complex GUI sequences into efficient programmatic calls. For example, this AgentS2 tool for spreadsheet manipulation replaces dozens of manual clicks with a single function:
def set_cell_values(self, cell_values: dict, app_name: str, sheet_name: str):
"""Set multiple cell values in a spreadsheet.
Args: cell_values: {"A2": "hello", "B3": 123.45}"""
return SET_CELL_VALUES_CMD.format(
cell_values=cell_values, app_name=app_name, sheet_name=sheet_name)
Automatic Scaling with Coding Agents. Inspired by CoACT-1 (Song et al., 2025), we adopt the multi-agent paradigm where an orchestrator dynamically delegates subtasks to either a GUI operator or a coding agent that executes Python/Bash scripts. This allows bypassing inefficient GUI sequences through direct programmatic execution. We extend this by mining the coding agent’s trajectories for reusable tools: when the coding agent solves subtasks programmatically, we employ an automatic LLM workflow to extract and refine these solutions into parameterized functions, with reflection steps and automated unit testing to ensure correctness. For example, from a trajectory where the coding agent modifies VS Code settings via script, we extract:
def add_vs_code_keybinding(key: str, command: str, when: str = ""):
"""Create or update a VS Code keybinding.
Args: key: "ctrl+j", command: "workbench.action.focusActiveEditorGroup"
Returns: {"path": "...", "action": "added", "backup": "..."}"""
2.2 Synthetic Data Engine for Hybrid Action Tasks
Large-scale synthetic training tasks for CUAs remain scarce, while existing resources are primarily test sets or complete trajectories with limited reproducibility. To address this, we developed a dual-pipeline synthetic data engine producing 17,000+ verifiable tasks for real-world computer use. Our engine operates through two complementary strategies: evaluator-first generation, ensuring verifiability, and instruction-first generation, creating contextually relevant tasks with diversity.
2.2.1 Evaluator-First Generation
This approach begins by collecting state-checking evaluators from computer environments—scripts that verify specific system states (e.g., file existence, application settings, UI elements). We use the atomic evaluator functions in OSWorld (Xie et al., 2024) to reprogram these evaluators by modifying parameters and compose multiple evaluators to create complex verification conditions. For example, combining a file-checker with a URL-checker validator creates a task requiring both file manipulation and browsing interaction.
Given these evaluator configurations, we prompt LLMs to generate corresponding tasks that would satisfy the verification conditions. For instance, the file-URL checker combination might generate tasks like “Navigate to the Python documentation page and download the PDF tutorial to your Documents folder,” which requires both web browsing to reach the correct URL and file system operations to verify the download. This ensures every generated task has a programmatic way to verify completion, critical for providing clear reward signals during RL training. This approach produces 4,000+ high-quality tasks with guaranteed verifiability.
2.2.2 Instruction-First Generation
Following Ramrakhya et al. (2025), this approach generates tasks based on observed system states. Agents explore computer environments through exploratory walks, reaching diverse UI states. At each state, we analyze the current interface and generate contextually appropriate tasks (e.g., “create a new spreadsheet” when in a file manager). Task completion is verified by an evaluator agent rather than predefined scripts, allowing flexibility in execution paths. This approach generates 12,000+ tasks that naturally arise from real usage patterns, complementing the systematic coverage of evaluator-first generation.
2.2.3 Workspace Simulation
A realistic workspace is crucial for generating meaningful tasks. When synthetic tasks require interaction with specific content, our pipeline triggers a content preparation workflow tailored to task requirements. For example, for code-related tasks, we fetch files from popular GitHub repositories—extracting Python scripts from Hugging Face repos or configuration files from trending projects. For image tasks, we retrieve open-source images from Wikipedia Commons matching relevant categories. For document editing, we generate synthetic documents via LLMs with task-appropriate content. This targeted approach ensures realistic task contexts: image editing tasks receive actual photos, code refactoring tasks get real implementations, and document tasks operate on properly formatted files. By matching content types to task requirements, we create scenarios that accurately reflect real-world computer use.
2.2.4 Complementary Design Rationale
In general, the two approaches serve distinct purposes. Evaluator-first generation produces complex, verifiable tasks ideal for RL training—code-based evaluators provide precise rewards without expensive trajectory verification. However, these tasks tend to be challenging due to evaluators’ design and multi-evaluator compositions. Instruction-first generation offers greater diversity through environment exploration, covering more real-world scenarios with naturally easier tasks. This complementary design ensures both reliable RL signals and broad task coverage. We further summarize detailed data statistics in Table 1.
Table 1: Comparison of our two synthetic data generation strategies.
Strategy Task Count Rollout SR (%) Avg. Difficulty Avg. Steps Total Samples Total Traj.
Evaluator-First 4K 29 Medium-Hard 6.8 33K 4.8K
Instruction-First 13K 45 Easy-Medium 6.5 149K 22K
2.3 Training a Foundation Agent with Hybrid Action
We train our foundation model using a two-stage approach: supervised fine-tuning on high-quality trajectory demonstrations followed by online reinforcement learning. This curriculum first establishes competency in hybrid action, then optimizes action selection between GUI primitives and programmatic interfaces through self-play.
2.3.1 Multi-Agent Rollout for Trajectory Generation
To generate high-quality training data, we deploy a multi-agent system comprising a Planner agent and a specialized Grounder agent. We use OpenAI o3 as the Planner, which operates in a ReAct framework (Yao et al., 2022) with Agent-S2-style prompting (Agashe et al., 2025) to enhance reasoning capabilities. The Planner strategically chooses between programmatic calls and GUI actions based on task context and available tools. When GUI interaction is needed, we employ GTA1-7B (Yang et al., 2025a) as the Grounding agent for precise visual localization, ensuring accurate element targeting in complex interfaces. For each synthetic task, we expose relevant programmatic tools to the Planner and perform 8 rollouts to capture diverse solution strategies. This process generates 26.8K successful trajectories demonstrating effective hybrid action strategies across our synthetic tasks.
2.3.2 Working Memory Mechanism
<memory>
Task: Create folder Favorites on bookmarks bar.
Progress: Chrome open, bookmarks bar visible.
Next: Access bookmark manager via Ctrl+Shift+O.
</memory>
Complex hybrid execution paths risk losing context as agents alternate between programmatic tools and GUI actions. We address this through an integrated working memory system using <memory></memory> tags, inspired by Bonatti et al. (2024). The agent autonomously maintains this memory—recording completed steps, extracted values, and intermediate results—ensuring coherent execution without external storage. The common memory content includes: (1) task objectives and constraints, (2) progress tracking across completed actions, and (3) information that must persist across steps (e.g., file paths, UI element states, intermediate values). For example, during a bookmark management task, the agent maintains structured state information as shown. This mechanism proves crucial for multi-step tasks requiring information persistence across action modality switches.
2.3.3 Stage 1: Supervised Fine-Tuning
We fine-tune multiple base models, including UI-TARS-1.5 (7B) (Qin et al., 2025) and OpenCUA (32B) (Wang et al., 2025b) on the 26.8K successful trajectories from the rollout system. To ensure balanced training across all trajectory steps, we create individual samples from each turn: for the $i$-th turn, we include messages up to that point but apply loss only to the $i$-th assistant response. This prevents overfitting to early trajectory steps while ensuring each action decision receives equal training weight, teaching the model proper hybrid action at every step of task execution.
2.3.4 Stage 2: Online Reinforcement Learning
While SFT provides behavioral foundations, mastering strategic action selection requires learning from exploration. The hybrid action space creates numerous solution paths for each task—some efficient, others suboptimal. Through online RL, agents can discover these optimal strategies via self-play.
We begin by filtering our evaluator-first tasks (4,000+) through 8 rollouts per task with the SFT model, identifying 1,000 tasks where the model succeeds at least once. We define task difficulty as the average success rate across these 8 rollouts. During training, we randomly sample tasks with difficulty scores in [0.4,0.8][0.4,0.8]—avoiding tasks that are too easy or too challenging to maximize learning efficiency within the model’s zone of proximal development.
For policy optimization, we employ a variant of GRPO (Shao et al., 2024) inspired by DAPO (Yu et al., 2025), with key modifications for our hybrid action setting. We remove KL regularization and implement a clip-higher strategy to encourage exploration of diverse action sequences.
To prevent regression toward GUI-only solutions, we design a reward function that incentivizes tool usage. The total reward for a trajectory $\tau$ is:
$R(\tau)=R_{\text{env}}(\tau)+R_{\text{tool}}(\tau)\,$, (1)
where $R_{\text{env}}(\tau)\in\{-1,1\}$ is the sparse environment reward (1 for task success, -1 for failure), and the tool-use reward is defined as:
$R_{\text{tool}}(\tau)=\begin{cases}0.3&\text{if }R_{\text{env}}(\tau)=1\text{ and }\tau\text{ contains tool calls},\\ 0&\text{otherwise}.\end{cases}$ (2)
This reward structure teaches the agent not just to succeed, but to succeed efficiently through strategic hybrid action. Notably, we exclude format rewards despite their common use in RL with LLMs. We found in empirical analysis that models struggle with complex tool syntax early in training, causing format penalties to dominate the learning signal and discourage outcome-based learning. By focusing solely on outcome and tool-use rewards, we enable the model to gradually master tool syntax through successful examples rather than punishment, leading to more robust learning. We propagate rewards to each action step and normalize by trajectory length for stable optimization.
3 Experiments
3.1 Experimental Setup
Table 2: Comparison of the state-of-the-art methods on the OSWorld benchmark. We split the results by steps and show the approach type in the second column. We report the success rate (%) as the evaluation metric in the fourth column. $\dagger$ denotes our reproduced results, averaged across 4 independent runs. Same-colored rows share the same base model.
Agent Method Model Category Open-Source Success Rate (%)
Max Steps: 15 Max Steps: 50
o3 (OpenAI, 2025) General Model 9.1 17.2
Claude 3.7 Sonnet (Anthropic, 2025) General Model 27.1 35.8
OpenAI CUA (OpenAI, 2025) Agentic Model 26.0 31.3
Jedi-7B w/ GPT-4o (Xie et al., 2025) Multi-Agent Framework 26.8 27.0
Agent S2 (Agashe et al., 2025) Multi-Agent Framework 27.0 34.5
Qwen2.5-VL-72B (Bai et al., 2025) General Model 4.4
UI-TARS-72B-DPO (Qin et al., 2025) Agentic Model 24.0 25.8
OpenCUA-7B (Wang et al., 2025b) Agentic Model 24.3 28.2
UI-TARS-1.5-7B (Qin et al., 2025) Agentic Model 24.5 27.3
OpenCUA-32B (Wang et al., 2025b) Agentic Model 29.7 34.1
UI-TARS-1.5-7B (Qin et al., 2025) Agentic Model 23.4 26.1
OpenCUA-32B (Wang et al., 2025b) Agentic Model 33.3 35.1
UltraCUA-7B-SFT Agentic Model 27.0 28.5
UltraCUA-7B-RL Agentic Model 28.9 30.2
UltraCUA-32B-SFT Agentic Model 39.0 41.5
UltraCUA-32B-RL Agentic Model 41.0 43.7
Benchmarks. We use OSWorld-Verified (Xie et al., 2024) as our primary benchmark. It is a realistic benchmark featuring a Ubuntu Desktop environment accessible through screen observations, comprising 369 tasks. OSWorld contains diverse tasks spanning common office suites, IDEs, and web browsers, designed to rigorously test an agent’s long-horizon planning and visual grounding abilities. Each task is self-contained with a deterministic starting state, a natural language goal, and an automated rule-based evaluator, ensuring reproducible and reliable assessment. To evaluate cross-platform generalization, we also test on WindowsAgentArena (Bonatti et al., 2024), which contains 154 real-world tasks in Windows 11 environments. This provides an out-of-domain evaluation since our models are primarily trained on Ubuntu-based tasks, testing the transferability of learned hybrid action strategies across operating systems.
Baselines. To demonstrate the effectiveness of our approach, we compare our final model against several strong baselines that isolate different components of the agent’s capabilities.
  • General Models: powerful, pre-trained vision-language models that are not specifically fine-tuned for GUI automation. We include leading models like Claude (Anthropic, 2025) and o3 (OpenAI, 2025) to establish a baseline for generalist, out-of-the-box performance.
  • Multi-Agent Frameworks: systems that orchestrate multiple components to solve computer-use tasks. These frameworks typically employ a planner-grounder architecture and may be enhanced with additional capabilities such as memory, experience replay, or the integration of a coding agent. Prominent examples include Agent-S2 (Agashe et al., 2025) and Jedi-7B (Xie et al., 2025).
  • Specialized Agentic Models: models that have been specifically fine-tuned or purpose-built for computer control and GUI-centric scenarios. This includes models like OpenAI CUA (OpenAI, 2025), UI-TARS (Qin et al., 2025) and OpenCUA (Wang et al., 2025b), which are trained on large datasets of computer interaction trajectories to specialize their abilities for this domain.
Training Details. Our models are fine-tuned for 3 epochs during the SFT stage with a learning rate of 2e-5. For the SFT stage, we sample 66K steps from trajectories with evaluator-first and instruction-first synthetic data, each 33K. The subsequent online RL stage is trained for 150 steps with a learning rate of 1e-6. All experiments are conducted on NVIDIA H100 GPUs. During training, we control the number of programmatic tools to limit the context length to 32K.
Evaluation Metrics. We use the following metrics to measure effectiveness and efficiency: 1) Success Rate (SR): Our primary metric. It is the percentage of tasks the agent successfully completes in a single attempt, as verified by the benchmark’s automated evaluators. 2) Pass@4: To account for the stochastic nature of LLM inference, we also report Pass@4. A task is marked as successful under this metric if the agent completes it correctly in at least one of four independent rollout attempts. 3) Trajectory Efficiency: We measure the number of steps an agent takes to complete a task successfully. Each step is either a GUI action or a programmatic tool call. A lower step count indicates higher efficiency.
Table 3: Detailed performance comparison of agent methods across different domains on OSWorld, measured by success rate (%) under the 15-step setting. Highlighted cells denote best results for each domain.
Methods Chrome Calc Impress Writer GIMP VSCode Multi Apps Thunderbird OS VLC Avg.
(46) (47) (47) (23) (25) (23) (101) (15) (24) (17)
(avg. runs)
UI-TARS-1.5-7B 32.5 8.4 22.8 36.9 48.5 43.5 7.1 38.3 34.1 27.0 23.4
(4.0) (4.0) (4.0) (4.0) (4.0) (4.0) (4.0) (4.0) (4.0) (4.0) (4.0)
UltraCUA-7B-SFT 40.7 13.3 26.9 46.7 57.7 34.8 8.4 41.1 36.0 30.3 27.0
(4.0) (4.0) (4.0) (4.0) (4.0) (4.0) (4.0) (4.0) (4.0) (4.0) (4.0)
UltraCUA-7B-RL 41.2 13.9 27.1 55.4 50.0 46.7 10.6 43.3 37.0 33.6 28.9
(4.0) (4.0) (4.0) (4.0) (4.0) (4.0) (4.0) (4.0) (4.0) (4.0) (4.0)
OpenCUA-32B 42.4 21.3 36.8 42.4 53.8 39.1 9.9 50.0 42.9 23.5 33.3
(4.0) (2.1) (2.2) (2.2) (4.0) (3.0) (1.9) (2.0) (1.8) (1.8) (2.5)
UltraCUA-32B-SFT 38.9 20.0 40.2 59.3 71.4 57.9 9.9 69.6 62.1 31.5 39.0
(4.0) (2.1) (2.2) (2.2) (4.0) (3.0) (1.9) (2.0) (1.8) (1.8) (2.5)
UltraCUA-32B-RL 40.6 25.7 40.2 62.5 70.0 54.3 14.9 72.1 64.0 33.3 41.0
(4.0) (2.1) (2.2) (2.2) (4.0) (3.0) (1.9) (2.0) (1.8) (1.8) (2.5)
3.2 Main Results
OSWorld Evaluation. Table 2 presents comprehensive results on the OSWorld benchmark across different step budgets. Our UltraCUA-7B achieves 28.9% success rate at 15 steps, surpassing all comparable 7B models, including the strong UI-TARS-1.5-7B baseline (23.4%) with a 23.5% relative improvement. More remarkably, UltraCUA-32B reaches 41.0% success rate, outperforming even closed-source systems like Claude 3.7 Sonnet (27.1%) and OpenAI CUA (26.0%). We also present the detailed evaluation results per domain in Table 3. Note that for OpenCUA-series models, due to the sub-optimal speed and infrastructure for inference, the overall average run is less than 4.
The results validate our hybrid action approach across model scales. While general-purpose models struggle without specialized training (e.g., Qwen2.5-VL-72B at 4.4% despite 72B parameters), our models achieve superior performance through strategic integration of programmatic tool calls. The consistent improvements from base models (UI-TARS-1.5-7B→UltraCUA-7B: +23.5%, OpenCUA-32B→UltraCUA-32B: +23.1%) demonstrate that hybrid action provides orthogonal benefits to agent capabilities.
Table 4: Out-of-domain evaluation on WindowsAgentArena.
Model SR (%)
Qwen2-VL-7B (w/ OpenCUA Data) 13.5
UI-TARS-1.5-7B 18.1
UltraCUA-7B 21.7
Cross-Platform Generalization. To assess generalization beyond the training domain, we evaluate on WindowsAgentArena without any Windows-specific fine-tuning. Table 4 shows that UltraCUA-7B achieves 21.7% success rate, outperforming both Qwen2-VL-7B trained with OpenCUA’s Windows data (13.5%) and UI-TARS-1.5-7B (18.1%). This 20% relative improvement over UI-TARS demonstrates that hybrid action strategies learned on Ubuntu effectively transfer to Windows environments, validating the domain-agnostic nature of our approach.
3.3 Ablation Studies
We conduct a series of ablation studies to dissect our framework and validate the contribution of its key components. These experiments isolate the impact of the hybrid action space, working memory, and reinforcement learning stage on agent performance.
3.3.1 The Impact of Hybrid Action
To validate the effectiveness of hybrid action, we examine its impact on both specialized agentic models and powerful multi-agent frameworks.
Impact on Specialized Models. We compare three configurations: (1) UI-TARS-1.5-7B (GUI-only baseline), (2) our model with tools disabled (UltraCUA-7B w/o Tools), and (3) our full model with hybrid action. Table 5 shows that hybrid action yields substantial improvements: success rate increases from 23.4% to 27.0% (+15.4% relative) while maintaining similar step counts. The addition of programmatic tools proves essential for effectiveness in complex automation tasks.
Impact on Multi-Agent Frameworks. To test whether hybrid action benefits extend to state-of-the-art systems, we evaluate our GTA1-7B + o3 rollout framework with and without programmatic tools. As shown in Table 5, hybrid action provides even larger gains in this setting: success rate improves from 44.0% to 48.2% (+9.5% relative) and average steps decrease by 14.9%. This demonstrates that hybrid action becomes increasingly valuable as the underlying system becomes more capable.
Table 5: Impact of hybrid action on different agent architectures. Hybrid action benefits both specialized models and multi-agent frameworks.
Model Configuration Success Rate (%) Pass@4 Avg. Steps
Agentic Models (Max Steps: 15)
UI-TARS-1.5-7B (GUI-Only) 23.4 33.3 9.31
UltraCUA-7B-SFT w/o Tools (GUI-Only) 25.1 34.3 9.24
UltraCUA-7B-SFT (Hybrid Action) 27.0 37.9 8.46
Commercial Models & Multi-Agent Framework (Max Steps: 50)
Claude-4-Sonnet 43.9
GTA1-7B + o3 w/o Tools 44.0 60.5 15.53
GTA1-7B + o3 (Hybrid Action) 48.2 62.4 13.22
3.3.2 The Importance of Reinforcement Learning
[…]
(a) Outcome Reward
(b) Format Reward
(c) Tool-call Pattern
Figure 3: Evolution of the agent’s behavior during reinforcement learning. Rewards are increasing as the number of failed tasks with tool-calls drops after RL.
3.3.3 Impact of Working Memory
We evaluate working memory by training models with and without <memory></memory> blocks in the SFT data, isolating the contribution of explicit state tracking. Table 6 shows consistent improvements from working memory: success rate increases from 25.4% to 27.0% (+6.3% relative) and average steps decrease slightly. While modest, these gains are meaningful for tasks requiring persistent state—file operations, form filling, and cross-application workflows. The efficiency improvement suggests memory helps agents avoid redundant actions like re-navigating to previously visited screens or re-extracting obtained information.
Table 6: Impact of working memory on model performance. Models are trained with identical data except for the presence of memory blocks.
Model Configuration Success Rate (%) Pass@4 Avg. Steps
UltraCUA-7B-SFT w/o Memory 25.4 37.1 8.56
UltraCUA-7B-SFT w/ Memory 27.0 37.9 8.46
Relative Improvement +6.3% +2.1% -1.2%
3.4 Analysis on Tool Use Patterns
To understand how our model leverages the hybrid action space, we analyze tool usage patterns across different application domains and task types.
Figure 4: Tool-call patterns across domains and models. Stronger models exhibit higher frequency and diversity.
Tool Usage Scales with Model Capability. Figure 4 reveals a clear correlation between model capability and tool usage sophistication. The multi-agent framework (GTA1-7B+o3) demonstrates extensive tool utilization with 60-80 calls and 8-10 unique tools per domain, while our single models show progressively conservative patterns—UltraCUA-32B uses tools moderately (20-40 calls) and UltraCUA-7B sparingly (0-20 calls). This pattern validates our hybrid action hypothesis: stronger models not only call tools more frequently but also leverage greater diversity, suggesting they better recognize when programmatic interfaces provide efficiency gains. The trend holds across all domains from office suites to development environments, confirming that effective hybrid action emerges naturally with increased model capability.
Table 7: OSWorld with OOD tools.
Configuration SR (%) Avg. Steps
UltraCUA-7B-SFT 27.0 8.46
w/ OOD tools 27.5 8.80
Out-of-Distribution Tool Generalization. We evaluate the model’s ability to utilize tools not seen during training by introducing new programmatic tools at inference time. These tools are unseen during training due to the context length limit. Table 7 shows that models can adapt to unseen tools, achieving modest performance gains (+1.9% relative SR). However, the increased steps suggest adaptation challenges—models may explore unfamiliar tools before selecting appropriate ones. This zero-shot tool generalization capability also extends beyond single-platform scenarios: Table 4 demonstrates that our model achieves 21.7% success rate on Windows tasks despite training exclusively on Ubuntu, outperforming baselines by leveraging its learned hybrid action strategies across platforms and tool ecosystems.
4 Conclusion
We introduced UltraCUA as foundation CUA models that bridge the critical gap between general-purpose GUI agents and specialized API-based agents. We achieve this through a novel hybrid action space that seamlessly integrates low-level GUI actions with high-level tool use. Our core contributions are a scalable pipeline for automated tool acquisition, a synthetic data engine for generating verifiable hybrid tasks, and a two-stage training curriculum to optimize strategic agent behavior. UltraCUA achieves state-of-the-art performance on the real-world benchmarks. Ablation studies confirm that the hybrid action space is the essential driver of this success, demonstrating a novel and more effective paradigm for building strong, robust and efficient agents for general computer control.
Acknowledgment
The authors would like to thank Harsh Agrawal, Dongxu Li, and Yutong Dai for valuable suggestions and feedback.
Apple and the Apple logo are trademarks of Apple Inc., registered in the U.S. and other countries and regions.
Appendix A Appendix
A.1 Related Work
Multimodal Agents for Computer Automation.
The ambition to create agents that can operate GUIs is long-standing, but has seen remarkable progress with the advent of Vision-Language Models (VLMs). Early approaches often relied on structured data like HTML or accessibility trees. More recent and generalizable agents operate directly from pixels and high-level instructions. In web automation, benchmarks like WebArena (Zhou et al., 2023) and Mind2Web (Deng et al., 2023) have driven the development of agents capable of complex online tasks. Similarly, in general computer control, works like CogAgent (Hong et al., 2023), Ferret UI You et al. (2024); Li et al. (2024) and OSWorld (Xie et al., 2024) have demonstrated agents that can navigate desktop environments, and AppAgent (Zhang et al., 2023) has shown similar capabilities on mobile devices. Current approaches to GUI automation can be broadly categorized into two paradigms. Multi-Agent systems employ specialized models for different subtasks—for instance, GPT-4o+Aria-UI (Yang et al., 2024) and GTA-1 (Yang et al., 2025a) combine a planner model with a dedicated grounder model, leveraging the strengths of each component for strategic planning and precise visual grounding, respectively. In contrast, Foundation Agent Models like UI-TARS (Qin et al., 2025), UI-TARS-2 (Wang et al., 2025a), OpenCUA (Wang et al., 2025b) and Ferret-UI Lite (Yang et al., 2025b) adopt an end-to-end approach, where a single unified model autonomously handles both planning and grounding tasks. While multi-agent systems benefit from modular design and specialized expertise, foundation models offer simpler deployment and potentially better coordination between planning and execution. A common thread among these powerful agents is their reliance on a primitive action space consisting of clicks, types, and scrolls. While this provides generality, it also leads to the brittleness and long-horizon planning challenges that our work directly addresses. Our contribution is the introduction of a hybrid action space that retains this generality while adding the efficiency and robustness of high-level tools.
Tool and API Augmentation for LLMs.
Parallel to the development of GUI agents, another line of research has focused on augmenting LLMs with the ability to use external tools and APIs. The seminal work of ToolFormer (Schick et al., 2023a) showed that models could learn to call APIs to access information they lack. This paradigm was rapidly scaled up by frameworks like ToolLLM (Qin et al., 2023b) and the Gorilla benchmark (Patil et al., 2023), which enabled models to select from thousands of real-world APIs. Furthermore, the concept of “tool-making” (Cai et al., 2023) has explored agents that can write their own tools when needed, a capability we incorporate into our tool acquisition pipeline. Recent advances have introduced reinforcement learning to tool-use training. ReTool (Feng et al., 2025) and ToolRL (Qian et al., 2025) pioneered the use of online RL for training end-to-end tool-use agents, demonstrating that reward signals alone can guide models to learn effective tool selection and usage strategies. These methods move beyond supervised learning on static datasets, allowing agents to discover optimal tool-use patterns through interaction and feedback. This RL-based paradigm aligns closely with our approach, where we employ online reinforcement learning to train agents that can strategically alternate between primitive GUI actions and high-level tool calls. While these tool-augmented agents are highly effective for structured, programmatic tasks, they typically operate in a non-visual, text-based environment and lack the ability to interact with the vast number of applications that do not expose an API. Our work bridges this gap, bringing the power of a rich tool ecosystem to the visually-grounded domain of GUI agents.
A.2 The Use of Large Language Models
We used large language models (LLMs) to assist with specific aspects of paper preparation. Specifically, LLMs were employed for: (1) language polishing and grammar checking to improve clarity and readability, (2) formatting suggestions to ensure compliance with conference style guidelines, and (3) recommendations for data visualization approaches to better present experimental results. All research ideas, experimental design, implementation, and core scientific contributions were developed by the authors without LLM assistance. The LLMs served purely as writing and presentation aids.
A.3 Details for Programmatic Tools
Table 8 summarizes the programmatic tools available across 10 different application domains on OSWorld. The collection comprises 881 tools in total, with individual domains offering between 4 (System) and 135 (VS Code) tools. These tools provide fine-grained control over desktop applications, enabling agents to perform tasks ranging from basic navigation (e.g., jump_to_next_tab) to complex application-specific operations (e.g., batch_spreadsheet_numeric_formatter). The comprehensive tool coverage ensures that agents can effectively automate diverse desktop workflows across different software environments.
Table 8: Overview of available tools across different domains.
Domain Tool Count Example Tools
Chrome 69 jump_to_next_tab, chrome_domain_data_wiper, open_downloads_page
GIMP 88 save_image_as, undo_last_action, swap_foreground_background_colors
LibreOffice General 41 open_find_and_replace, open_print_preview, open_hyperlink_dialog
LibreOffice Calc 114 spreadsheet_column_formula_injector, batch_spreadsheet_numeric_formatter
LibreOffice Impress 75 set_line_spacing_1, insert_non_breaking_space, apply_subscript
LibreOffice Writer 123 select_to_start_of_next_page, select_to_start_of_paragraph
System 4 open_system_terminal_and_execute, open_app_or_filename
Thunderbird 119 open_message_in_conversation, delete_message_permanently
VLC 83 set_video_as_wallpaper, volume_up, jump_1_minute_forward
VS Code 135 add_vs_code_keybinding, vscode_exclude_folders, search_within_current_file
Total 881
A.4 Details for Synthetic Tasks
We generated a comprehensive synthetic dataset of 17,864 tasks across 10 application domains using two complementary approaches. As shown in Table 9, the evaluator-first approach contributed 4,387 high-quality tasks with complex multi-step instructions, while the instruction-first approach generated 13,477 tasks to ensure broad coverage of application functionalities.
The dataset spans diverse applications from productivity tools (LibreOffice suite with 5,885 combined tasks) to specialized software like GIMP (1,121 tasks) and development environments like VS Code (1,990 tasks). Chrome represents the largest single-domain category with 2,826 tasks, reflecting the importance of web interactions. The multi-apps category (2,113 tasks) specifically tests cross-application workflows. Task complexity varies from simple operations (e.g., “Change the text alignment to Center”) to sophisticated procedures requiring multiple coordinated actions (e.g., “Convert video to MP4 and save with a new filename”), ensuring comprehensive evaluation of agents’ GUI navigation and task execution capabilities.
Table 9: Overview of synthetic data generation across different domains.
Domain Eval-First Instr-First Example Instructions Total
Chrome 751 2,075 “Find hotels in Paris for 2 adults for three nights starting next Friday and sort the list by lowest price.” “Restore the previous session pages in Google Chrome.” 2,826
GIMP 401 720 “Please replace the current white backdrop with a solid green color, but keep the black circle in the centre exactly as it is.” “In GIMP, navigate to the Display section and set the check style to Medium checks.” 1,121
LibreOffice Calc 651 1,496 “Open the spreadsheet and make the entire header row (row 1) bold.” “Protect the sheet Sheet2 in LibreOffice Calc.” 2,147
LibreOffice Impress 501 1,397 “Make every slide in this deck use a solid dark-green background (RGB 0 128 0). I’d like all the pages to share that exact colour so the presentation looks consistent.” “Add a video from /videos/video3.mov to slide 3 in LibreOffice Impress.” 1,898
LibreOffice Writer 851 989 “Change the default font in LibreOffice Writer to Calibri.” “Change the text alignment to Center in LibreOffice Writer.” 1,840
OS/System 301 1,197 “I accidentally created a file called ’draft.txt’ on my Desktop. Please delete it completely so it’s no longer there.” “View the partitioning table of the disk named {disk_name} in the Disks app.” 1,498
Thunderbird 351 1,084 “Create a new folder named ’ToSort’ inside the Local Folders section.” “Import contacts from Windows Mail into Thunderbird.” 1,435
VLC 330 666 “Open the cat photo in VLC and set it as my desktop wallpaper.” “Play the current video in VLC Media Player.” 996
VS Code 250 1,740 “Could you open VS Code and create a new text file named ’meeting_notes.txt’ inside the folder ’/home/user/notes’? Make sure to save the file before you finish.” “Search for the term Data Structure in the document and highlight it in LibreOffice Writer” 1,990
Multi-apps 2,113 “Change the desktop wallpaper to Desert on the Ubuntu desktop.” “Search for JavaScript in Brave settings and enable it.” 2,113
Total 4,387 13,477 17,864
A.5 Qualitative Examples
To illustrate the practical advantages of our hybrid action paradigm, we present three representative examples in Figures 5, 6, and 7. These cases highlight how UltraCUA strategically selects between high-level programmatic tools and low-level GUI actions to enhance efficiency, tackle complex problems, and ensure robust execution.
First, the email-starring task (Figure 5) exemplifies the agent’s capacity for intelligent and fluid alternation between control modes. The process begins with a precise low-level GUI click to select the target “Bills” folder, effectively setting the context. Immediately following this, the agent switches to high-level, reliable tool calls—select_all and add_or_remove_star—to execute the core bulk operation. This strategic handoff from a specific GUI action to general-purpose tools ensures both precision and operational robustness.
In the second example (Figure 6), the agent is asked to clear YouTube browsing history. Instead of relying on a potentially brittle sequence of clicks through menus, it initiates the workflow with a single programmatic tool call, open_history_page, to navigate directly to the correct settings page. Subsequently, it seamlessly transitions to primitive GUI actions—typing into a search field and clicking buttons—to perform the more nuanced task of filtering and deleting the specific entries. This demonstrates a practical fusion of programmatic speed for navigation and GUI flexibility for manipulation.
Finally, a more complex scenario in Figure 7 showcases the model’s ability to automate workflows that are intractable for purely GUI-based agents. When tasked with batch-processing images on the desktop, UltraCUA correctly identifies the need for a scripted solution. It programmatically opens a system terminal, installs the necessary software (imagemagick), and proceeds to write and execute a multi-line shell script to automate the entire process. This ability to generate and utilize code represents a significant leap in problem-solving capability.
Figure 5: An example of UltraCUA-32B helping processing emails. The agent alternates between low-level actions and programmatic tool calls smartly, leading to efficient completion of the task.
Figure 6: An example of UltraCUA-32B helping clear certain Chrome history with hybrid action. The agent calls prothe grammatic tool at the first step to assist in going into the desired page directly.
Figure 7: An example of UltraCUA-32B helping processing images with hybrid action. The model starts coding at the very first step by calling the terminal tool, and finally writes a bash script and executes it to make the task successful.